# GigaSLAM: Large-Scale Monocular SLAM with Hierachical Gaussian Splats

### v1

@@ Abstract
Tracking and mapping in large-scale, unbounded outdoor environments using only monocular RGB input presents substantial challenges for existing SLAM systems. Traditional Neural Radiance Fields (NeRF) and 3D Gaussian Splatting (3DGS) SLAM methods are typically limited to small, bounded indoor settings. To overcome these challenges, we introduce GigaSLAM [^1] https://github.com/DengKaiCQ/GigaSLAM, the first NeRF/3DGS-based SLAM framework for kilometer-scale outdoor environments, as demonstrated on the KITTI and KITTI 360 datasets. Our approach employs a hierarchical sparse voxel map representation, where Gaussians are decoded by neural networks at multiple levels of detail. This design enables efficient, scalable mapping and high-fidelity viewpoint rendering across expansive, unbounded scenes. For front-end tracking, GigaSLAM utilizes a metric depth model combined with epipolar geometry and PnP algorithms to accurately estimate poses, while incorporating a Bag-of-Words-based loop closure mechanism to maintain robust alignment over long trajectories. Consequently, GigaSLAM delivers high-precision tracking and visually faithful rendering on urban outdoor benchmarks, establishing a robust SLAM solution for large-scale, long-term scenarios, and significantly extending the applicability of Gaussian Splatting SLAM systems to unbounded outdoor environments.
@@ 1 Introduction
Simultaneous localization and mapping (SLAM) from a single monocular video is a longstanding challenge in computer vision [3]. Conventional SLAM approaches [25, 24, 4, 7] seek accurate camera pose tracking and high-quality map geometry. Recent advances in neural radiance fields (NeRF) [22] and 3D gaussian splatting (3DGS) [15] have inspired the SLAM community [46] to enhance map capabilities—enabling rich appearance encoding and realistic rendering from free viewpoints. Novel SLAM techniques using NeRFs and 3DGS have shown great promise for online mapping and rendering [57, 58, 55, 14, 21]. These capabilities open new possibilities for applications in AR/VR, autonomous driving, and drone navigation by allowing users and agents to render from different perspectives.
@@
Despite these advancements, one key challenge remains for such systems: current NeRF and 3DGS-based SLAM frameworks are often limited in scene scale, with most relying on RGB-D depth priors for accurate mapping and tracking in larger scenes [57, 14, 21]. This is due to two main reasons: the limitations of scene representations and challenges in global alignment. Implicit methods, such as NeRFs [57, 58, 55], have limited representational capacity and are often confined to bounded regions. Explicit methods, like Gaussian splatting [14, 21], are not memory-efficient, as their size scales with scene growth, impacting computational and memory efficiency. The second challenge lies in global alignment. In large-scale scenes, loop closure is crucial, as it effectively reduces global drift. However, most NeRF and 3DGS SLAM methods rely on incremental gradient-based registration, which is not well-suited for global alignment under such scenarios. Additionally, it remains unclear how to deform and update the map after loop closure without compromising rendering performance. As a result, existing NeRF and 3DGS SLAM methods are restricted to small-scale, bounded indoor environments and often depend on RGB-D data for robust scene alignment.
@@
To address these challenges, we present GigaSLAM, a Gaussian Splatting-based SLAM framework designed to scale to large, outdoor, long-term, unbounded environments. The core of GigaSLAM’s technical contribution is a novel hierarchical sparse voxel representation designed for large-scale rendering-capable SLAM, with each level encoding Gaussians at varying levels of detail. This map representation has two key advantages for scaling: 1) it is boundless, dynamically expanding as the camera moves; 2) it enables content-aware, efficient rendering at large scale through Level-of-Details (LoD) representation, reducing the need to load hundreds of millions of Gaussians for a single frame. We further enhance map geometry and camera pose tracking accuracy using a data-driven monocular metric depth module [27]. Finally, we integrate Bag-of-Words loop closure detection [8] and design a comprehensive post-closure Splats map update to maintain accuracy in rendering. GigaSLAM has been shown to scale up to tens of kilometers of travel distance in urban driving scenarios [10, 16]. It achieves robust and accurate tracking with the ability to render from any viewpoint, all on a single GPU. To our knowledge, no existing framework offers this capability.
@@
We validate our methods on large-scale outdoor scenes from the KITTI and KITTI 360 dataset [10, 16]. Our results indicate that this approach is significantly more accurate and robust, outperforming other monocular SLAM methods in average tracking performance on long-term outdoor datasets. Additionally, existing NeRF and 3D Gaussian Splatting-based SLAM frameworks struggle to handle these large-scale, unbounded scenarios effectively. For the indoor dataset, our approach achieves tracking and rendering performance comparable to state-of-the-art 3D GS SLAM methods [21] tailored for indoor scenes with monocular RGB images. Our contributions are as follows: 1) GigaSLAM, a novel Gaussian Splats-based SLAM framework for large-scale, unbounded environments; 2) a hierarchical map representation for dynamic growth and level-of-detail rendering in large-scale SLAM; 3) an efficient loop closure procedure applicable to Gaussian splats map representations.
@@ 2 Related Work
Traditional SLAM
Modern SLAM systems generally treat localization and mapping as a joint optimization problem [3], aiming to estimate a robot’s state (position and orientation) from video inputs. ORB-SLAM [25] (and its updated versions, ORB-SLAM2 [24] and ORB-SLAM3 [4]) is a well-known example, utilizing feature points and keyframes for monocular SLAM. In recent SLAM research, deep learning models are often used to solve specific sub-problems within SLAM and are then embedded in traditional optimization-based pipelines. $\nabla$ SLAM [13] leverages automatic differentiation to unify gradient learning with SLAM in a way that treats the SLAM system as a differentiable computation graph. Similarly, CodeSLAM [2] applies an autoencoder-based approach to achieve dense monocular SLAM by encoding geometric information of images. Additionally, some works [41, 17] have embedded Bundle Adjustment into end-to-end differentiable networks, with DROID-SLAM [42] being a leading example that integrates Dense Bundle Adjustment into direct optical flow estimation.
NeRF-based SLAM
The emergence of Neural Radiance Fields (NeRF) [22] has sparked the development of several high-performance SLAM algorithms. iMAP [39] was the first to introduce NeRF-based techniques for high-fidelity scene reconstruction in SLAM. NICE-SLAM [57] leverages a grid-based multi-level local scene representation, enabling it to handle complex indoor environments. NICER-SLAM [58] introduces geometric and optical flow constraints and a simple warping loss to enhance geometric consistency, allowing for monocular input. Point-SLAM [32], built on Point-NeRF [50], uses point-based neural fields for explicit scene representation, but requires RGB-D input and lacks integration with traditional SLAM techniques. GO-SLAM [55], building on NICE-SLAM, introduces loop closure detection to ensure global consistency in 3D reconstruction. However, since NeRF requires predefined scene boundaries, which are difficult to determine in outdoor environments, most algorithms become inapplicable in such scenarios.
Gaussian Splatting-based SLAM
The introduction of 3D Gaussian Splatting (3DGS) [15] has opened up new possibilities for SLAM. [14, 21] were the first to apply 3DGS to SLAM. SplaTAM [14] uilizing an online tracking and mapping pipeline, Gaussian primitives, and differentiable contour-guided optimization for both mapping and localization. MonoGS [21] uses the explicit nature of Gaussian functions, introducing geometric validation and regularization to handle ambiguities in incremental dense 3D reconstruction, successfully achieving indoor scene localization and reconstruction with RGB-D or RGB input. But RGB input of MonoGS leads to a relatively worse tracking and mapping quality than the RGB-D input. GS-SLAM [51] proposes an adaptive extension strategy to efficiently reconstruct new scene geometries and improve previous map regions. RGBD GS-ICP SLAM [12] integrates Generalized Iterative Closest Point (G-ICP) with 3DGS, enhancing localization accuracy. Splat-SLAM [33] is a method built upon the tracking module of [42] and achieves the most accurate performance in indoor scenarios with RGB-only input. But it has not yet been tested in outdoor environments. Overall, most 3DGS-based SLAM methods still rely on RGB-D input and are primarily focused on indoor scenarios with few tests on long-term outdoor datasets.
@@ 3.4 Loop Closure
where $S_{i}$ is the absolute similarity of keyframe $i$, the first term is the smoothness term between consecutive keyframes, the second term represents the error for the loop closure between keyframes $j$ and $k$, $\Delta S$ is the relative similarity between keyframes. Details could be found in the supplementary material. Our approach applies a rigid transformation to all voxels across different levels to maintain spatial consistency. This ensures that anchor points remain correctly aligned across the updated camera coordinate frames. Subsequently, a re-voxelization process is required to adaptively refine the map structure, which includes deduplication step. To efficiently manage redundant voxels introduced by loop closure corrections, we leverage the spatial hashing mechanism (Eq. 11), which enables fast lookup and merging of overlapping voxels. This ensures a compact and consistent voxel representation while preserving rendering quality after the update, making our approach particularly well-suited for hierarchical voxel-based SLAM.
(a) Sequence 07 & 08
(b) Sequence 09 & 10
Figure 4: Rendering results on the KITTI dataset are shown for MonoGS (with RGB input) Splat-SLAM and our proposed method. Our method maintains stable rendering across extended outdoor sequences, while MonoGS and Splat-SLAM struggle on this scene. Due to space limitations, rendering comparisons for other sequences can be found in the suppl.
@@
4 Experiments
4.1 Experimental Setup
4.1.1 Dataset and Metrics
We evaluate our system on two datasets: KITTI [10], KITTI 360 [16]. KITTI is the primary dataset due to its kilometer-scale, long sequences, offering a challenging outdoor environment for SLAM. Unlike other methods, which are limited to smaller datasets or indoor scenes, our approach effectively handles large-scale outdoor scenarios.
For tracking accuracy, we report the Absolute Trajectory Error (ATE) [38] of the keyframes. Map quality is evaluated using photometric rendering metrics such as Peak Signal-to-Noise Ratio (PSNR) [49], Structural Similarity Index Measure (SSIM) [49], and Learned Perceptual Image Patch Similarity (LPIPS) [54], which collectively capture both pixel-level and perceptual differences in rendered images.
4.1.2 Implementation Details
4.2 KITTI Dataset
Figure 5: Trajectory estimation of different SLAM methods on sequence 00 of the KITTI dataset with RGB input. Our method demonstrates stable tracking over long outdoor scenes, unlike the scale drift in DROID-SLAM and the tracking failure of MonoGS.
Table 1 presents the tracking performance of our method on the KITTI [10] dataset (Data from DPV-SLAM [19]). Overall, our method demonstrates strong tracking capabilities, particularly in long and complex sequences. Figure 5 further illustrates a comparison of camera trajectories among different SLAM methods on the KITTI dataset’s 00 sequence. Our method demonstrates stable tracking performance in outdoor, long-sequence scenarios, maintaining consistent and accurate camera pose estimates. By contrast, DROID-SLAM [42] shows substantial scale drift when processing long-distance data, leading to inferior performance on extended sequences. This limitation indicates that DROID-SLAM lacks robustness for long-duration tracking tasks. MonoGS [21], however, encounters even greater challenges on the KITTI dataset: after processing a few hundred frames, it has crashed and was unable to continue tracking. Furthermore, as shown in Figure 4 and Table 2, the error estimated camera poses in MonoGS severely impact both mapping and tracking accuracy. This prevents MonoGS from performing effective tracking and mapping on the long, outdoor sequences typical of the KITTI dataset. In more detail, our method achieves stable and accurate camera pose estimates across long-distance, outdoor sequences in the KITTI dataset, demonstrating robustness that existing methods lack.
4.3 KITTI 360 Dataset
(a) Seq. 0002 (Ours)
(b) Seq. 0004 (Ours)
(c) Seq. 0006 (Ours)
(d) Seq. 0005 (Ours)
(e) Seq. 0005 (DROID)
(f) Seq. 0006 (DROID)

### v2

# GigaSLAM: Large-Scale Monocular SLAM with Hierarchical Gaussian Splats

@@ Abstract
Tracking and mapping in large-scale, unbounded outdoor environments using only monocular RGB input presents substantial challenges for existing SLAM systems. Traditional Neural Radiance Fields (NeRF) and 3D Gaussian Splatting (3DGS) SLAM methods are typically limited to small, bounded indoor settings. To overcome these challenges, we introduce GigaSLAM [^1] https://github.com/DengKaiCQ/GigaSLAM, the first RGB NeRF / 3DGS-based SLAM framework for kilometer-scale outdoor environments, as demonstrated on the KITTI, KITTI 360, 4 Seasons and A2D2 datasets. Our approach employs a hierarchical sparse voxel map representation, where Gaussians are decoded by neural networks at multiple levels of detail. This design enables efficient, scalable mapping and high-fidelity viewpoint rendering across expansive, unbounded scenes. For front-end tracking, GigaSLAM utilizes a metric depth model combined with epipolar geometry and PnP algorithms to accurately estimate poses, while incorporating a Bag-of-Words-based loop closure mechanism to maintain robust alignment over long trajectories. Consequently, GigaSLAM delivers high-precision tracking and visually faithful rendering on urban outdoor benchmarks, establishing a robust SLAM solution for large-scale, long-term scenarios, and significantly extending the applicability of Gaussian Splatting SLAM systems to unbounded outdoor environments.
@@ 1 Introduction
Simultaneous localization and mapping (SLAM) from a single monocular video is a longstanding challenge in computer vision [3]. Conventional SLAM approaches [27, 26, 4, 8] seek accurate camera pose tracking and high-quality map geometry. Recent advances in neural radiance fields (NeRF) [24] and 3D gaussian splatting (3DGS) [17] have inspired the SLAM community [50] to enhance map capabilities—enabling rich appearance encoding and realistic rendering from free viewpoints. Novel SLAM techniques using NeRFs and 3DGS have shown great promise for online mapping and rendering [63, 64, 61, 16, 23]. These capabilities open new possibilities for applications in AR/VR, autonomous driving, and drone navigation by allowing users and agents to render from different perspectives. Unlike traditional 3D GS or NeRF methods that rely on a pre-reconstructed static scene from SfM[36], SLAM-based approaches must dynamically build and update the 3D scene online. This online nature introduces unique challenges, especially when handling loop closures and global scene corrections.
@@
Despite these advancements, one key challenge remains for such systems: current NeRF and 3DGS-based SLAM frameworks are often limited in scene scale, with most relying on RGB-D depth priors for accurate mapping and tracking in larger scenes [63, 16, 23]. This is due to two main reasons: the limitations of scene representations and challenges in global alignment. Implicit methods, such as NeRFs [63, 64, 61], have limited representational capacity and are often confined to bounded regions. Their reliance on manually pre-defined scene bounding boxes becomes impractical in expansive outdoor environments with dynamic scales and undefined boundaries. Furthermore, the prevalent dense volumetric grid representations exhibit cubic space complexity $O(n^{3})$, which incurs prohibitive memory and computational costs when scaling to outdoor scenes spanning thousands of cubic meters. Explicit methods, like Gaussian splatting [16, 23], are not memory-efficient, as their size scales with scene growth, impacting computational and memory efficiency. Meanwhile, concurrent work like OpenGS-SLAM [58] integrates 3R modules [53] with 3D GS, achieving capability on short hundred-meter-scale outdoor sequences (Waymo[43]). However, its Transformer-based 3R design incurs $O(n^{2})$ memory costs due to self-attention, limiting scalability to kilometer-scale trajectories. Current 3R-based forward 3D reconstruction frameworks remain limited in scalability, as no prior method has demonstrated robust performance on sequences exceeding kilometer-level spatial extents. The second challenge lies in global alignment. In large-scale scenes, loop closure is crucial, as it effectively reduces global drift. However, most NeRF and 3DGS SLAM methods rely on incremental gradient-based registration, which is not well-suited for global alignment under such scenarios. As a result, existing NeRF and 3DGS SLAM methods are restricted to small-scale, bounded indoor environments or short-term outdoor scenes and often depend on RGB-D data for robust scene alignment.
@@
To address these challenges, we present GigaSLAM, a Gaussian Splatting-based SLAM framework designed to scale to large, outdoor, long-term, unbounded environments. The core of GigaSLAM’s technical contribution is a novel hierarchical sparse voxel representation designed for large-scale rendering-capable SLAM, with each level encoding Gaussians at varying levels of detail. This map representation has two key advantages for scaling: 1) it is boundless, dynamically expanding as the camera moves; 2) it enables content-aware, efficient rendering at large scale through Level-of-Details (LoD) representation, reducing the need to load hundreds of millions of Gaussians for a single frame. We further enhance map geometry and camera pose tracking accuracy using a data-driven monocular metric depth module [29]. Finally, we integrate Bag-of-Words loop closure detection [9] and design a comprehensive post-closure Splats map update to maintain accuracy in rendering. GigaSLAM has been shown to scale up to tens of kilometers of travel distance in urban driving scenarios [11, 18, 55, 12]. It achieves robust and accurate tracking with the ability to render from any viewpoint, all on a single GPU. To our knowledge, no existing framework offers this capability.
@@
We validate our methods on large-scale outdoor scenes from the KITTI, KITTI 360, 4 Seasons and A2D2 dataset [11, 18, 55, 12]. Our results indicate that this approach is significantly more accurate and robust, outperforming other monocular SLAM methods in average tracking performance on long-term outdoor datasets comparing to current 3D GS SLAM methods [23, 35] which are tailored for indoor scenes with monocular RGB images. Therefore, existing NeRF and 3D Gaussian Splatting-based SLAM frameworks struggle to handle these large-scale, unbounded scenarios effectively. Our contributions are as follows: 1) GigaSLAM, a novel Gaussian Splats-based SLAM framework for large-scale, unbounded environments; 2) a hierarchical map representation for dynamic growth and level-of-detail rendering in large-scale SLAM; 3) an efficient loop closure procedure applicable to Gaussian splats map representations.
@@ 2 Related Work
Traditional SLAM
Modern SLAM systems are typically formulated as joint optimization problems [3], where the goal is to estimate a robot’s pose (position and orientation) from video input. A representative system is ORB-SLAM [27], along with its extended versions ORB-SLAM2 [26] and ORB-SLAM3 [4], which use feature points and keyframes for monocular SLAM. In more recent studies, researchers have incorporated deep learning models to solve specific sub-modules within SLAM, which are then integrated into traditional optimization-based frameworks. For example, the authors of $\nabla$ SLAM [15] propose using automatic differentiation to model SLAM as a differentiable computation graph. Similarly, CodeSLAM [2] introduces an autoencoder-based representation for dense monocular SLAM by learning compact geometric descriptors from images. Several works [45, 19] have also embedded Bundle Adjustment into end-to-end differentiable networks. DROID-SLAM [46] exemplifies this trend by integrating Dense Bundle Adjustment directly into an optical flow estimation pipeline.
NeRF-based SLAM
The introduction of Neural Radiance Fields (NeRF) [24] has inspired a range of SLAM systems that leverage neural rendering for mapping and localization. iMAP [42] is one of the first works to explore NeRF-based scene reconstruction for SLAM. Building on this, NICE-SLAM [63] adopts a grid-based hierarchical representation to improve efficiency in indoor mapping. NICER-SLAM [64] further introduces geometric and optical flow constraints, along with a warping loss, to improve consistency while supporting monocular input. The authors of Point-SLAM [34] employ point-based neural fields for explicit scene representation, though their method requires RGB-D input and does not integrate with traditional SLAM architectures. GO-SLAM [61] extends NICE-SLAM by incorporating loop closure detection to enhance global reconstruction consistency. However, since NeRF-based systems generally require pre-defined scene boundaries, they struggle to scale to unbounded outdoor environments, limiting their applicability in such scenarios.
Gaussian Splatting-based SLAM
The emergence of 3D Gaussian Splatting (3DGS) [17] has motivated researchers to explore its potential for SLAM. The earliest SLAM adaptations of 3DGS include SplaTAM [16] and MonoGS [23]. In SplaTAM, the authors propose an online SLAM framework that employs Gaussian primitives and differentiable contour-guided optimization for both tracking and mapping. MonoGS takes advantage of the explicit and compact nature of Gaussians, introducing geometric validation and regularization to address ambiguities in dense reconstruction; the method supports both RGB and RGB-D inputs, though RGB-only input leads to reduced tracking and mapping quality. GS-SLAM [56] introduces an adaptive extension strategy for efficient map updates and reconstruction of novel regions. RGBD GS-ICP SLAM [14] incorporates Generalized ICP with 3DGS to improve localization precision. Splat-SLAM [35], based on the tracking framework of DROID-SLAM [46], achieves state-of-the-art accuracy in indoor scenes using only RGB input, but its performance in outdoor environments remains unverified. VPGS-SLAM [7] is an independent work shortly after our study. It maintains LiDAR sensors as input, thereby incurring higher hardware costs typical of such LiDAR-based systems. Overall, current 3DGS-based SLAM systems tend to rely on RGB-D or LiDAR data and are mostly evaluated on indoor scenes, with limited validation on large-scale or long-term outdoor sequences.
@@ 3.4 Loop Closure
where $S_{i}$ is the absolute similarity of keyframe $i$, the first term is the smoothness term between consecutive keyframes, the second term represents the error for the loop closure between keyframes $j$ and $k$, $\Delta S$ is the relative similarity between keyframes. Details could be found in the supplementary material.
After pose update, Our approach applies a rigid transformation to all voxels across different levels to maintain spatial consistency. We update all voxels in the global coordinate frame to align with the optimized camera poses. Let $\mathbf{p}_{i}\in\mathbb{R}^{3}$ be the $i$-th anchor point, originally associated with the $j$-th camera pose. Given the original pose $\mathbf{T}_{\text{old}}^{(j)}\in\text{SE}(3)$ and the optimized pose $\mathbf{T}_{\text{new}}^{(j)}\in\text{SE}(3)$, the updated position of the anchor point is given by:
$\mathbf{p}_{i}^{\text{new}}=\mathbf{T}_{\text{new}}^{(j)}\cdot\left(\mathbf{T}_{\text{old}}^{(j)}\right)^{-1}\cdot\begin{bmatrix}\mathbf{p}_{i}\\ 1\end{bmatrix}.$ (13)
To reduce memory consumption for large-scale voxels, we process the update in batches. The entire anchor point set of voxels $\mathcal{P}=\{\mathbf{p}_{1},\dots,\mathbf{p}_{N}\}$ is divided into $M$ disjoint subsets $\mathcal{B}_{1},\dots,\mathcal{B}_{M}$, each of size $B\ll N$. Within each batch $\mathcal{B}_{m}$, the update rule remains consistent, but applied locally:
$\forall\,\mathbf{p}\in\mathcal{B}_{m},\quad\mathbf{p}^{\text{new}}=\mathbf{T}_{\text{new}}^{(\pi(\mathbf{p}))}\cdot\left(\mathbf{T}_{\text{old}}^{(\pi(\mathbf{p}))}\right)^{-1}\cdot\begin{bmatrix}\mathbf{p}\\ 1\end{bmatrix},$ (14)
where $\pi(\mathbf{p})$ maps each point $\mathbf{p}$ to its associated camera pose index. This ensures that anchor points remain correctly aligned across the updated camera coordinate frames. Subsequently, a re-voxelization process (Eq. 5) is required to adaptively refine the map structure. To efficiently manage voxels introduced by loop closure corrections, we leverage the spatial hashing mechanism (Eq. 11), which enables fast lookup of updated voxels.
4 Experiments
Figure 4: This figure evaluates rendering and geometric quality via global view on KITTI 06 Sequence. Splat SLAM exhibits severe scale inconsistency , whereas our method maintains precise scale coherence even in long outdoor sequences.
@@ 4 Experiments
4.1 Experimental Setup
4.1.1 Dataset and Metrics
We evaluate our system on datasets: KITTI [11], KITTI 360 [18] 4 Seasons[55], A2D2[12]. KITTI is the primary dataset due to its kilometer-scale, long sequences, offering a challenging outdoor environment for SLAM. Unlike other methods, which are limited to smaller datasets or indoor scenes, our approach effectively handles large-scale outdoor scenarios.
For tracking accuracy, we report the Absolute Trajectory Error (ATE) [41] of the keyframes on three dataset: KITTI [11], KITTI 360 [18] 4 Seasons[55]. Due to the absence of GT pose matrices in the A2D2 dataset, ATE computation becomes infeasible. We thus visualize our algorithm’s tracking performance through projecting of its trajectory on Google Map in Figure 8. Mapping quality is evaluated using photometric rendering metrics such as Peak Signal-to-Noise Ratio (PSNR) [54], Structural Similarity Index Measure (SSIM) [54], and Learned Perceptual Image Patch Similarity (LPIPS) [60], which collectively capture both pixel-level and perceptual differences in rendered images.
4.1.2 Implementation Details
4.2 KITTI Dataset
(a) Sequence 07 & 08
(b) Sequence 09 & 10
Figure 6: Rendering results on the KITTI dataset are shown for MonoGS (with RGB input) Splat-SLAM and our proposed method. Our method maintains stable rendering across extended outdoor sequences, while MonoGS and Splat-SLAM struggle on this scene.
Table 1 presents the tracking performance of our method on the KITTI [11] dataset (data from DPV-SLAM [21]). Overall, our method shows strong tracking performance, particularly on long and complex sequences. Figure 7 compares camera trajectories on sequence 00, where our approach maintains stable and accurate pose estimates. In contrast, DROID-SLAM [46] exhibits significant scale drift on long sequences, indicating limited robustness. MonoGS [23] performs worse—it crashes after a few hundred frames and fails to continue tracking. As shown in Figure 6 and Table 3, the inaccurate poses from MonoGS degrade both mapping and tracking quality, making it unsuitable for KITTI’s long, outdoor sequences. Our method, however, delivers consistently accurate poses over extended trajectories, showing robustness unmatched by these baselines.
Figure 4 qualitatively compares reconstructions from our method (Ours LoD 0), MonoGS, and Splat SLAM on KITTI 06. From left to right, it shows global reconstruction, ground truth pointcloud and trajectory, and zoom-ins on local geometry. MonoGS outputs sparse, noisy maps with major detail loss. Splat SLAM captures more structure but suffers from scale inconsistency due to failed loop closure. Our method reconstructs more consistent geometry throughout the sequence and better preserves scene details, even in distant areas. Zoom-in views highlight our strength in capturing thin structures and avoiding over-splatting.
Figure 7: (Left) Trajectory estimation of different SLAM methods on sequence 00 of the KITTI dataset with RGB input. Our method demonstrates stable tracking over long outdoor scenes, unlike the scale drift in DROID-SLAM and the tracking failure of MonoGS. (Right) Our method on 4 Seasons dataset.
KITTI 360 Dataset 4 Seasons Dataset
Methods Avg. $0$ 0002 0003 0004 0005 0006 0007 0009 0010 Avg. Business Campus Office Loop Old Town Neighborhood City Loop
seq. frames 8497 11518 14607 1031 11587 6743 9699 3396 14056 3836 20200 17280 15177 28999 11121 28424
seq. length (m) 6971.088 8403.15 11501.28 1378.73 9975.28 4690.74 7979.92 4887.51 10579.68 3343.49 4906.55 3132.38 3710.66 5258.60 2078.37 10352.74
contains loop - -
DROID-SLAM [46] 193.307 110.07 233.87 10.79 169.11 139.02 113.81 577.39 165.34 220.36 / OOM 175.63 OOM 158.19 OOM
Ours (w/o LC) 61.402 25.53 56.59 29.13 116.15 20.51 28.52 203.02 16.83 56.35 99.098 8.81 36.85 69.71 7.48 372.64
Ours (w/ LC) 47.107 17.72 34.98 27.55 40.54 20.46 19.12 197.33 13.22 53.03 92.950 7.33 21.12 67.89 6.28 362.13
Table 2: Camera Tracking Results (ATE RMSE [m] $\downarrow$) on the KITTI 360 & 4 Seasons Dataset. OOM stands for Out-of-Memory. marks a closed-loop sequence with just one hardly detectable start-end loop point; DBoW failed to identify this latent loop closure.
@@ 4.2 KITTI Dataset
4.3 KITTI 360, 4 Seasons & A2D2 Dataset